# WindVE: Collaborative CPU-NPU Vector Embedding

### v2

@@ 4.1. System Design
4.2. Queue Manager
4.2.1. Overall Procedure
@@ 12                         $cpu\_queue\_push(query_{k})$
4.2.2. Linear Regression Based Queue Depths Estimation
@@
To address this challenge, WindVE is equipped with a fast estimator designed to determine the upper limits of concurrency for both CPUs and NPUs/GPUs. The methodology is grounded in the observed linear relationship between the processing latency and the concurrency, as mentioned by SLSC (Cheng et al., 2024) and Mooncake (Qin et al., 2024). Therefore, the relationship between latency $t_{proc,d}^{C_{d}}$ and concurrency $C_{d}$, $d\in\{NPU,CPU\}$ can be expressed as follows:
(12) $t_{proc,d}^{C_{d}}=\alpha_{d}\cdot C_{d}+\beta_{d},d\in\{NPU,CPU\}$
@@
4.2.3. Theoretical Effectiveness Analysis
@@
(13) $t_{proc}$ $=t_{comp}+t_{io}+t_{model}$
@@
(14) $\alpha_{CPU}$ $>\alpha_{NPU}$
(15) $\beta_{CPU}$ $>\beta_{NPU}$
@@
(16) $\beta_{CPU}$ $>\beta_{NPU}$
(17) $t-\beta_{NPU}$ $>t-\beta_{CPU}$
(18) $\alpha_{NPU}\cdot C_{NPU}$ $>\alpha_{CPU}\cdot C_{CPU}$
(19) $\frac{\alpha_{NPU}}{\alpha_{CPU}}$ $>\frac{C_{CPU}}{C_{NPU}}$
As detailed in Section 3.2, there is a direct correlation between the ratio $\frac{C_{CPU}}{C_{NPU}}$and the extent of extended concurrency. Inequality 19 delineates that the upper limit of this concurrency is governed by the ratio $\frac{\alpha_{NPU}}{\alpha_{CPU}}$. Consequently, the narrower the performance disparity between NPUs/GPUs and CPUs, the greater the concurrency advantages that WindVE can leverage.
@@
(20) $\frac{\alpha_{NPU}}{\alpha_{CPU}}>\frac{C_{CPU}}{C_{NPU}}$
(21) $\frac{\frac{\Delta t}{\alpha_{CPU}}}{\frac{\Delta t}{\alpha_{NPU}}}>\frac{C_{CPU}}{C_{NPU}}$
(22) $\frac{\Delta C_{CPU}}{\Delta C_{NPU}}>\frac{C_{CPU}}{C_{NPU}}$
(23) $\frac{C_{CPU}+\Delta C_{CPU}}{C_{NPU}+\Delta C_{NPU}}>\frac{C_{CPU}}{C_{NPU}}$
@@
Input: $NPU_{i},CPU_{j},\forall i\in[1,I],j\in[1,J]$
Output: $device\_main,device\_auxiliary,worker\_num\_main$,
$worker\_num\_auxiliary,heter\_enable$
1 procedure $DeviceDetect(NPU_{i},CPU_{j})$
[…]
4 $worker\_num\_main=0$
5 $worker\_num\_auxiliary=0$
6 if npu is available then
7       if heterogeneous computing is enabled then
8             /* heter_enable=true */
[…]
11             $worker\_num\_main=I$
12             $worker\_num\_auxiliary=J$
13
14      else
[…]
17             $worker\_num\_main=J$
18             $worker\_num\_auxiliary=0$
19
20else
[…]
23       $worker\_num\_main=J$
24       $worker\_num\_auxiliary=0$
25       $heter\_enable=false$
26
27return $device\_main,device\_auxiliary,heter\_enable$
Algorithm 2 Device detector logic of CPU and NPU
4.3. Device Detector
@@
The device detector logic is shown in Algorithm 2.
4.4. CPU Affinity and Numa Settings for ARM CPUs
@@
5. Experiments
In this section, a series of experiments for overall performance and the system scalability have been delivered on the CPU-NPU collaborative system. Experiment settings are introduced in Section 5.1. These settings serve as defaults in experiments for overall performance demonstrated in Section 5.2 and 5.4 for system scalability if not mentioned specifically.
@@ 5. Experiments
FlagEmbedding concurrency 44 96 84 172
WindVE concurrency 44 + 8 96 + 22 84 + 1 172 + 8
concurrency improvement 18.2% 22.3% 1.2% 4.7%
Table 1. Overall performance of WindVE against FlagEmbedding ob bge model under 1s and 2s limit.
@@
PyTorch concurrency 48 112 128 256
WindVE concurrency 48 + 11 112 + 30 128 + 6 256 + 20
concurrency improvement 22.9% 26.7% 4.6% 7.8%
Table 2. Overall performance of WindVE against PyTorch on jina model under 1s and 2s limit.
Figure 4. Fitting curves and stress tests on (A) Tesla V100 GPU, (B) two Intel Xeon E5 2690 CPUs, (C) one Atlas 300I DUO NPU, (D) and two Kunpeng 920 CPUs.
device Tesla V100 Intel Xeon E5 Atlas 300I DUO Kunpeng 920
queue depth under 1s limit linear regression 40 8 84 2
@@
queue depth under 2s limit linear regression 96 20 195 15
@@
Table 3. Queue Depth Prediction through linear regression and through stress tests.
Figure 5. Scalability experiment with varied input query length in Tesla V100 and Intel Xeon E5 2690 under 1s and 2s limit.
5.1. Experiment Settings
5.1.1. Testbed
To validate the CPU-NPU collaborative system, experiments have been run in CPUs with both x86 and ARM architectures, and in both NPUs and GPUs. Two type of machines have been employed in experiments: one is a GPU machine with Tesla V100 GPUs (Nvidia, 2024) and Intel Xeon E5-2680 CPUs (Intel, 2024) with x86 architecture and 48 cores in total; another is Atlas 800 Inference Server (Model: 3000) (Huawei, 2024b) with Atlas 300I DUO NPUs (Huawei, 2024a) and Kunpeng 920 CPUs (Huawei, 2024c) with ARM architecture and 128 cores in total.
5.1.2. Embedding Models
The mainly embedding model used in the experiments is bge-large-zh-v1.5 model (Xiao et al., 2023). Bge-large-zh-v1.5 is an embedding model to convert input queries to a output vector with 1024 fp32 elements, with around 326 million parameters. Besides, jina model ((Günther et al., 2023), 570M parameters and 8192 output length) is also adopted as supplementary.
5.1.3. Workloads
5.1.4. Baselines
The embedding framework used in experiments is FlagEmbedding (of Artificial Intelligence, 2024), which is based on PyTorch. FlagEmbedding is an embedding framework using BGE (BAAI General Embedding), a toolkit focused on retrieval-augmented LLMs with a set of projects in inference, fine-tuning, evaluation, datasets, tutorials, and research. As for jina embedding model, it can be executed directly by the transformer module of PyTorch.
5.1.5. Experiment Procedures
Figure 6. Scalability experiment with varied CPU core numbers in Tesla V100 and Intel Xeon E5 2690 under 1s and 2s limit.
5.2. Overall Performance
Initially, we employ the queue depth estimator detailed in Section 4.2.2 to estimate the possible maximum concurrency for NPUs/GPUs and CPUs. Next, the queue depth is fine-tuned according to the estimated value with CPUs and NPUs/GPUs running collaboratively. The overall performance of WindVE aganist counterparts on bge and jina models are shown in Table 1 and Table 2, respectively.
Comparison on bge model. From the results, the best concurrency improvement of 22.3% is given by the collaboration of one Tesla V100 GPU and two Intel Xeon E5 2690. Based on the analysis given in 3.2, this improvement can reduce 18.6% deployment cost when using maximum concurrent query numbers to deploy, and save up to 22.3% deployment cost when using average query numbers.
Comparison on jina model. Apparently, both CPU and NPU/GPU exhibit more concurrency on jina than bge, e.g., the concurrency of Kunpeng 920 is 20 when the time limit is 2s, much higher than the value (8) on bge. The concurrency improvement reaches 26.7% with Intel Xeon E5 2690, compared to the baseline that uses Tesla V100 only. Thus, the deployment cost of jina model can be reduced by 21.1% and 26.7% on maximum and average cases, respectively.
Three notable phenomena emerge from Table 1:
@@ 5.2. Overall Performance
Section 4.2.3 has already furnished the theoretical underpinnings for the existence of these phenomena, and this section delivers experimental evidence to validate the analysis.
@@
5.3. Queue Depth Estimator Evaluation
An individual assessment of the queue depth estimator has been conducted, and the results are presented in Table 3, showcasing the queue depth predictions produced from linear regression, stress tests with the increment step of 8, and the final fine-tuned queue depth achieved during CPU-NPU collaboration. The results indicate that the predictions yielded by linear regression demonstrate either superior or comparable efficacy to those obtained through stress tests, with the exception of the cases involving the Atlas 300I DUO and Kunpeng 920 under the 2-second limit. The comparatively better performance of the linear regression based prediction is attributed to an overly large increment step in the stress tests, which resulted in missing the peak of maximum concurrency. Nevertheless, Atlas 300I DUO and Kunpeng 920 generate a larger number of outliers when mapping the relationship between concurrency and end-to-end latency. This results in a less accurate prediction of queue depths when compared to the results of stress testing.
5.4. Scalability Experiments
@@
Scalability with CPU cores. Figure 6 shows that fewer CPU core numbers worsen the CPU concurrency on Intel Xeon E5 2690. With 1s limit, using less than 44 CPU cores does not bring any performance benefit, because the loss of computing ability leads to the dramatical increase of CPU latency. This boundary can be reduced to 36 given 2s limit, as these CPU cores iare enough to finish the embedding within 2s. In practical, with 2 CPUs with 128 cores (usually organized as 4 numas), we can utilize at most 96 cores (corresponding to the latter 3 numas) because our main program as well as the system programs runs on the first numa. However, even with more CPUs equipped in the server, the concurrency can not be improved continuously after a border, due to the bottleneck of host memory bandwidth (only tens of Gigabytes per second).
@@
6. Conclusions
In this paper, we found out the relations between embedding service deployment costs, the throughput and the system maximum concurrency. To solve the overtime issue of deployment using average processing ability, we crafted a queue manager to fully utilize idle CPUs in servers to enlarge system concurrency without sacrificing processing speed. To be more specific, the queue manager offloads excessive queries to CPUs to avoid overtimes in NPUs/GPUs. Additionally, we provide theoretical guidance on the calibration of its pivotal parameter for maximize concurrency. With the queue manager as the core component, a flexible and scalable system framework WindVE for vector embedding has been developed. Unlike other works focused on reducing embedding latency, WindVE aims to uplift throughput through concurrency increase. Mathematical expressions have also been summarized to allow the performance benefit estimation of using the CPU offlaoding and to guide key hyperparameter searching. Finally, we evaluated the performance of WindVE through experiments, and proved that the concurrency can be enlarged up to 22.3% when running in one Tesla V100 GPU and two Intel Xeon E5 2690 CPUs. Scalability experiments have also been conducted to show the impact of input query length and used CPU cores. We summarized that WindVE can effectively boost system concurrency with short input queries, with as many cores as possible, and with small CPU-to-NPU performance gap.

### v4

@@ IV-A System Design
IV-B Queue Manager
IV-B1 Overall Procedure
@@ 12                         $cpu\_queue\_push(query_{k})$
IV-B2 Linear Regression Based Queue Depths Estimation
@@
To address this challenge, WindVE is equipped with a fast estimator designed to determine the upper limits of concurrency for both CPUs and NPUs/GPUs. The methodology is grounded in the observed linear relationship between the processing latency and the concurrency, as mentioned by SLSC [45] and Mooncake [46]. Therefore, the relationship between latency $t_{proc,d}^{C_{d}}$ and concurrency $C_{d}$, $d\in\{NPU,CPU\}$ can be expressed as follows:
$t_{proc,d}^{C_{d}}=\alpha_{d}\cdot C_{d}+\beta_{d},d\in\{NPU,CPU\}$ (12)
@@
IV-B3 Theoretical Effectiveness Analysis
@@
$t_{proc}$ $=t_{comp}+t_{io}+t_{model}$ (13)
@@
$\alpha_{CPU}$ $>\alpha_{NPU}$ (14)
$\beta_{CPU}$ $>\beta_{NPU}$ (15)
@@
$\beta_{CPU}$ $>\beta_{NPU}$ (16)
$t-\beta_{NPU}$ $>t-\beta_{CPU}$ (17)
$\alpha_{NPU}\cdot C_{NPU}$ $>\alpha_{CPU}\cdot C_{CPU}$ (18)
$\frac{\alpha_{NPU}}{\alpha_{CPU}}$ $>\frac{C_{CPU}}{C_{NPU}}$ (19)
As detailed in Section III-B, there is a direct correlation between the ratio $\frac{C_{CPU}}{C_{NPU}}$and the extent of extended concurrency. Inequality 19 delineates that the upper limit of this concurrency is governed by the ratio $\frac{\alpha_{NPU}}{\alpha_{CPU}}$. Consequently, the narrower the performance disparity between NPUs/GPUs and CPUs, the greater the concurrency advantages that WindVE can leverage.
@@
$\frac{\alpha_{NPU}}{\alpha_{CPU}}>\frac{C_{CPU}}{C_{NPU}}$ (20)
$\frac{\frac{\Delta t}{\alpha_{CPU}}}{\frac{\Delta t}{\alpha_{NPU}}}>\frac{C_{CPU}}{C_{NPU}}$ (21)
$\frac{\Delta C_{CPU}}{\Delta C_{NPU}}>\frac{C_{CPU}}{C_{NPU}}$ (22)
$\frac{C_{CPU}+\Delta C_{CPU}}{C_{NPU}+\Delta C_{NPU}}>\frac{C_{CPU}}{C_{NPU}}$ (23)
@@
IV-C Device Detector
@@
IV-D CPU Affinity and Numa Settings for ARM CPUs
@@
V Experiments
In this section, a series of experiments for overall performance and the system scalability have been delivered on the CPU-NPU collaborative system. Experiment settings are introduced in Section V-A. These settings serve as defaults in experiments for overall performance demonstrated in Section V-B and V-D for system scalability if not mentioned specifically.
V-A Experiment Settings
V-A1 Testbed
To validate the CPU-NPU collaborative system, experiments have been run in CPUs with both x86 and ARM architectures, and in both NPUs and GPUs. Two type of machines have been employed in experiments: one is a GPU machine with Tesla V100 GPUs [47] and Intel Xeon E5-2680 CPUs [48] with x86 architecture and 48 cores in total; another is Atlas 800 Inference Server (Model: 3000) [49] with Atlas 300I DUO NPUs [50] and Kunpeng 920 CPUs [51] with ARM architecture and 128 cores in total.
V-A2 Embedding Models
The mainly embedding model used in the experiments is bge-large-zh-v1.5 model [43]. Bge-large-zh-v1.5 is an embedding model to convert input queries to a output vector with 1024 fp32 elements, with around 326 million parameters. Besides, jina model ([44], 570M parameters and 8192 output length) is also adopted as supplementary.
V-A3 Workloads
V-A4 Baselines
The embedding framework used in experiments is FlagEmbedding [32], which is based on PyTorch. FlagEmbedding is an embedding framework using BGE (BAAI General Embedding), a toolkit focused on retrieval-augmented LLMs with a set of projects in inference, fine-tuning, evaluation, datasets, tutorials, and research. As for jina embedding model, it can be executed directly by the transformer module of PyTorch.
V-A5 Implementations
To utilize the CPU processing without slowing down the tasks in NPUs, we reserved 1/4 of the total cores for Arm-based Kunpeng CPUs with low indexes for NPU processing. Additionally, we ensured the NPU’s maximum concurrency when utilizing CPU offloading to avoid sacrificing NPU processing.
V-A6 Experiment Procedures
@@
FlagEmbedding conc. 44 96 84 172
WindVE conc. 44 + 8 96 + 22 84 + 1 172 + 8
conc. improvement 18.2% 22.3% 1.2% 4.7%
TABLE I: Overall performance of WindVE against FlagEmbedding ob bge model under 1s and 2s limit.
@@
PyTorch conc. 48 112 128 256
WindVE conc. 48 + 11 112 + 30 128 + 6 256 + 20
conc. improvement 22.9% 26.7% 4.6% 7.8%
TABLE II: Overall performance of WindVE against PyTorch on jina model under 1s and 2s limit.
Figure 4: Fitting curves and stress tests on (A) Tesla V100 GPU, (B) two Intel Xeon E5 2690 CPUs, (C) one Atlas 300I DUO NPU, (D) and two Kunpeng 920 CPUs.
device queue depth Tesla V100 Intel Xeon E5 Atlas 300I DUO Kunpeng 920
1s limit linear regression 40 8 84 2
@@
2s limit linear regression 96 20 195 15
@@
TABLE III: Queue Depth Prediction through linear regression and through stress tests.
Figure 5: Scalability experiment with varied input query length in Tesla V100 and Intel Xeon E5 2690 under 1s and 2s limit.
Figure 6: Scalability experiment with varied CPU core numbers in Tesla V100 and Intel Xeon E5 2690 under 1s and 2s limit.
V-B Overall Performance
Initially, we employ the queue depth estimator detailed in Section IV-B2 to estimate the possible maximum concurrency for NPUs/GPUs and CPUs. Next, the queue depth is fine-tuned according to the estimated value with CPUs and NPUs/GPUs running collaboratively. The overall performance of WindVE against counterparts on bge and jina models are shown in Table I and Table II, respectively.
Comparison on bge model. From the results, the best concurrency improvement of 22.3% is given by the collaboration of one Tesla V100 GPU and two Intel Xeon E5 2690 with a less than 0.1 ms overhead. Besides, by applying CPU offloading, the average CPU utilization rises from 8% to 80%. Based on the analysis given in III-B, this improvement can reduce 18.6% deployment cost when using maximum concurrent query numbers to deploy, and save up to 22.3% deployment cost when using average query numbers.
Comparison on jina model. Apparently, both CPU and NPU/GPU exhibit more concurrency on jina than bge, e.g., the concurrency of Kunpeng 920 is 20 when the time limit is 2s, much higher than the value (8) on bge. The overhead stays less than 0.1ms, and the average CPU utilization rises from 9% to 81%. The concurrency improvement reaches 26.7% with Intel Xeon E5 2690, compared to the baseline that uses Tesla V100 only. Thus, the deployment cost of jina model can be reduced by 21.1% and 26.7% on maximum and average cases, respectively.
Three notable phenomena emerge from Table I:
@@ V-B Overall Performance
Section IV-B3 has already furnished the theoretical underpinnings for the existence of these phenomena, and this section delivers experimental evidence to validate the analysis.
@@
V-C Queue Depth Estimator Evaluation
An individual assessment of the queue depth estimator has been conducted, and the results are presented in Table III, showcasing the queue depth predictions produced from linear regression, stress tests with the increment step of 8, and the final fine-tuned queue depth achieved during CPU-NPU collaboration. The results indicate that the predictions yielded by linear regression demonstrate either superior or comparable efficacy to those obtained through stress tests, with the exception of the cases involving the Atlas 300I DUO and Kunpeng 920 under the 2-second limit. The comparatively better performance of the linear regression based prediction is attributed to an overly large increment step in the stress tests, which resulted in missing the peak of maximum concurrency. Nevertheless, Atlas 300I DUO and Kunpeng 920 generate a larger number of outliers when mapping the relationship between concurrency and end-to-end latency. This results in a less accurate prediction of queue depths when compared to the results of stress testing.
V-D Scalability Experiments
@@
Scalability with CPU cores. Figure 6 shows that fewer CPU core numbers worsen the CPU concurrency on Intel Xeon E5 2690. With 1s limit, using less than 44 CPU cores does not bring any performance benefit, because the loss of computing ability leads to a surge of CPU latency. This boundary can be reduced to 36 given 2s limit, as these CPU cores are enough to finish the embedding within 2s. In practical, with 2 CPUs with 128 cores (usually organized as 4 numas), we can utilize at most 96 cores (corresponding to the latter 3 numas) because our main program as well as the system programs runs on the first numa. However, even with more CPUs equipped in the server, the concurrency can not be improved continuously after a border, due to the bottleneck of host memory bandwidth (only tens of Gigabytes per second).
@@
VI Conclusions
In this paper, we identifies the relationship between embedding service costs, throughput, and concurrency. To address deployment delays caused by average processing times, we developed a queue manager that offloads excess queries to idle CPUs, boosting system concurrency without slowing processing. Theoretical guidance helps calibrate its key parameter for maximum concurrency. This manager forms the core of WindVE, a system designed to increase throughput via concurrency, unlike others focusing on latency. Mathematical expressions allow performance estimation and hyperparameter tuning. Experiments show WindVE boosts concurrency by up to 22.3% using one GPU and two CPUs. Scalability tests indicate WindVE is most effective with short queries, ample CPU cores, and minimal CPU-GPU performance gaps.